OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 + 8x + 18

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

(8/2)^2=16 y=x^2+8x+16-16+18 y=(x+4)^2+2

OpenStudy (anonymous):

THANKYOU